# Learning Successor Features with Distributed Hebbian Temporal Memory

### v2

@@ 5 Conclusion
This paper introduces a novel probabilistic Factorial-HMM-like algorithm, DHTM, for learning an observation sequence model in non-stationary environments that uses local Hebbian-like learning rules. DHTM is scalable to multiple feature variables by employing sparse distributed representations and sparse factor function implementation using segments, which inspire biologically plausible multicomponent neural models. In contrast to methods that use Monte-Carlo trajectory sampling for future state probability estimation, our method can perform belief propagation, so each prediction step adds a constant amount of computation, allowing a long planning horizon.
Experiments show that our memory model can quickly learn and unlearn observation sequences and transition dynamics, even in changing environments. In that case, methods that rely on backpropagation, like LSTM and CSCG, are difficult to train and require rigorous trajectory buffer management, a common problem in RL tasks.
One of the main limitations of the DHTM is that its temporal context is completely random, as it is formed on the fly using only the local information available. As a result, the mechanism of context formation allows only limited generalizations between different trajectories. However, this is what makes fast learning possible. Combining this fast learning with more slow learning that allows generalizations is a promising direction. Forming Successor Features combined with a two-level hierarchy of DHTM layers may provide the next step to a more generalized sequence model. The first (fast) DHTM layer may be used to form Successor Features both for planning and as current state representation, which is then observed by the second (slow) DHTM layer. We hypothesize that because SFs should provide a better generalization than the raw features, the second layer should be able to form a more versatile model of the environment.
6 Impact Statements
This paper presents research that aims to advance the field of machine learning. There are numerous potential societal consequences of our work, none of which we believe require special emphasis here.
Appendix A Encoding and Decoding Observations
Because our model is designed to work with sparse distributed representations and the testing environments do not provide observations as SDRs by default, an encoding procedure is required. For this task, we use a modified version of the Spatial Pooler (SP) [Cui et al., 2017, Mnatzaganian et al., 2017], a distributed noise-tolerant online clustering neural network algorithm that converts input binary patterns into SDRs with fixed sparsity while retaining pairwise similarity [Kuderov et al., 2023]. The SP algorithm learns a spatial specialization of neurons’ receptive fields using the local Hebbian rule and k-WTA ($k$ winners take all) inhibition [Oster et al., 2009]. Here we outline the main differences from the “vanilla” version of the SP algorithm described in Cui et al. [2017].
During an agent’s decision-making process pipeline, the SP encoder accepts a current observation $o$ and transforms it to a latent state SDR $z$. Observation $o$ should be an array of binary states. To transform an RGB image into a binary array, we perform the following processing steps, simulating a simple event-based camera:
  1. 1.
    Transform image to greyscale.
  2. 2.
    Calculate difference between current and previous greyscale image $d$.
  3. 3.
    Determine average difference over the entire image $\tau=\mathrm{mean}(d)$.
  4. 4.
    For pixels that have $d$ above $\tau$ set $1$, for others set 00.
In terms of processing, our SP encoder functions as a standard artificial neural network with a k-WTA binary activation function.:
$\text{overlaps}_{i}$ $=\beta_{i}W_{i}o$ (14)
$z_{i}$ $=\mathbb{I}\left[i\in\text{kWTA}(\text{overlaps})\right],$ (15)
where $o$—a binary observation vector, $W_{i}$—a row-vector representing $i$-th neuron’s connection weights (where non-existing connections have zero weights), $\text{overlaps}_{i}$—a value representing the strength of the input pattern recognition with the neuron $i$  [^1] While the name “overlap” does not exactly reflect its meaning in our SP modification, because it is not a binary overlap between a receptive field and an input pattern, we kept it on purpose to refer to the similar term commonly used for the original SP., $\beta_{i}$—an $i$-th neuron boosting value, $z_{i}$—an $i$-th bit of an output SDR, $\mathbb{I}[\dots]$—an indicator function, kWTA—a $k$-winners-take-all activation function returning $k$ indices of the neurons with the highest overlap.
One difference between the “vanilla” SP algorithm and ours is that we do not distinguish between potential and active neural connections. Because all [existing] connections are active, they all participate in calculating overlaps. In the overlaps calculation, non-binary, that is, real-valued weights are used, similar to artificial neural networks, as shown in equation 14. Furthermore, each neuron has a fixed capacity to produce neurotransmitters, which it distributes between its synaptic connections. This means that we keep all neuron weights normalized and summed to one. While it achieves the same Hebbian learning with homeostatic plasticity as the original SP, the exact formula is slightly different:
$\tilde{W}_{i}$ $=W_{i}+\alpha z_{i}\frac{\text{RF}_{i}\odot o}{\sum_{j}{\text{RF}_{i}\odot o}}$
$W_{i}$ $\leftarrow\frac{\tilde{W}_{i}}{\sum_{j}{\tilde{W}_{j}}},$ (16)
where $\tilde{W}_{i}$—a row of new $i$-th neuron weights before normalization, $\alpha$—learning rate, $z_{i}$—a binary value representing the current activity state of the $i$-th neuron, $RF_{i}$—an $i$-th row of the binary connectivity matrix representing an $i$-th neuron receptive field, $\odot$—elementwise product, $o$—a binary observation vector.
The original SP algorithm has several drawbacks, including encoding instability caused by an innate homeostatic plasticity mechanism known as boosting, which helps neurons specialize and increases overall adaptability but makes memorization tasks more difficult, and slow processing on large inputs such as images, where an encoding overhead becomes noticeable when compared to overall model timings around 1k input size.
The introduction of the newborn stage, which follows the ideas proposed in Dobric et al. [2022], solves an encoding instability problem. The newborn stage of a spatial pooler occurs during the early stages of its learning process, when its neurons are expected to specialize. The boosting, which is intended to aid in the specialization process, is activated only during the newborn stage and its scale gradually decreases from the configured value to zero. Boosting remains turned off during an encoder’s “adulthood”, reducing the possibility of spontaneous re-specialization.
To reduce processing overhead, we use a much more sparsified connection matrix than in the original SP version. We randomly initialize connections with 40-60% sparsity, which is typical for the “vanilla” SP. Then, during the newborn stage, we gradually prune the vast majority of the weakest connections, resulting in neurons that are highly specialized due to their small receptive fields. We typically configure the final receptive field size in relation to the average input pattern size (usually 25-200% of it, resulting in 0.1-10% connections sparsity). For example, if binary input patterns have on average 100 active bits out of 1000, we can set the target size of receptive fields to 25, which is 25% of the active input size and corresponds to 2.5% connection matrix sparsity. As a result, the spatial pooler’s instability (and thus adaptiveness!) becomes even more limited in the adult stage.
Because of its soft discretization (from the distributed representation) and clusterization properties, we expect SP to assist the model with input sequence memorization and an environment transition dynamics generalization tasks in addition to the encoding itself. However, because the SP encoder learns online, particularly during the newborn stage, its output representation can be highly unstable during the early stages, potentially resulting in a performance drop.
To visualize and debug an encoded observation, we also learn a decoder, which is a linear neural layer learned locally with gradient descend on the MSE error between the predicted reconstruction and the actual observation.
Appendix B Value Function Decomposition
@@
In our agent model, we approximate the reward function R(s) as a sum: $R_{t}=\frac{1}{n}\sum^{n}_{k=1}{r(\varphi^{k}_{t})\varphi^{k}_{t}}$, where $r(\varphi^{k}_{t})$ is a reward associated with state $\varphi^{k}_{t}$, $n$–number of feature variables. Then, similarly to the Successor Representation idea (see Section 2.3), the value function can be represented as:
@@
$=\frac{1}{n}\sum^{n}_{k=1}\sum_{j}\sum^{\infty}_{l=0}\gamma^{l}p(\varphi^{k}_{t+l+1}=j\ |\ h^{k}_{t})r^{k}_{j}$
$=\frac{1}{n}\sum^{n}_{k=1}\sum_{j}M^{k}_{j}(h^{k}_{t})r^{k}_{j},$ (17)
where $M^{k}_{j}(h^{k}_{t})=\sum^{\infty}_{l=0}\gamma^{l}p(\varphi^{k}_{t+l+1}=j\ |\ h^{k}_{t})$, $h_{t}=(h^{1}_{t},...,h^{n}_{t})$—hidden state vector of variables $\{H^{k}_{t}\}_{k}$.
Appendix C Biological Interpretation
@@
The neuronal activity at timestep $t$ can be thought to carry messages $M^{k}_{t-1}$. Messages are propagated through synapses of dendritic segments, which correspond to factors $F^{k}_{c}$. Dendritic segments of biological neurons are known to be coincidence detectors of its synaptic input [Stuart and Spruston, 2015]. We use the notion of dendritic segment to sparsely represent context factors $F^{k}_{c}$, as each factor value corresponds to a particular combination of states (or active cells).
Feature variables $\Phi^{k}_{t}$ may be considered to represent cells of a granular layer (L4), as they are known to be the main hub for sensory excitation for L2/3. L2/3 cells that have common sensory input from the layer L4 are modelled as columns for particular feature states $\mathrm{col(\varphi^{k}_{t})}$ [Mountcastle, 1997].
@@ Appendix D Implementation details
As mentioned in Section 3.3, we incorporate DHTM as a part of an RL agent, which has a memory model and a feature reward function. All tested memory models share the same pipeline—they learn sequences of encoded binary observations (i.e. SDRs that we get from Spatial Pooler encoder described in Appendix A) concatenated with one-hot encoded actions.
@@ D.1 LSTM
LSTM baseline was implemented with a single LSTMCell from PyTorch library [Paszke et al., 2019]. It is supported by an additional symexp-layer to encode input before passing it to the LSTM cell and a symexp-layer to decode the LSTM cell’s output from the LSTM’s hidden state back to the input representation, where symexp activation function, $\mathrm{symexp}(x)=\mathrm{sign}(x)e^{\lvert x\rvert-1}$, is a reverse of symlog function: $\mathrm{symlog}=\mathrm{sign}(x)\log{(\lvert x\rvert+1})$.
For the Gridworld task, LSTM with a hidden state size of 100 is trained on a buffer of 100 trajectories with a decreasing learning rate starting at 0.02. 25 trajectories are left out for validation on each epoch. The training is terminated when the validation loss is not decreasing more than 3 epochs in a row. After that, 30% of the buffer is replaced with new trajectories, and training continues.
We also incorporated some notion of random variables in hidden states by splitting the hidden state of the tested LSTM into groups. In all experiments the hidden state represents $10$ categorical variables with $10$ states. That is, LSTM is forced to learn $10$ categorical distributions with multi-cross-entropy loss to explain the observed sequences, which is a somewhat close to the multi-categorical hidden state representation used in DreamerV2/V3 [Hafner et al., 2023]. The idea of using symexp activation function, mentioned above, is inspired by Dreamer too, and is used to remedy the problem of learning extreme probability values. Without symexp the neural network has to represent zero probability with high negative logit values and one-probability with high positive logit values, which is hard to reach with low learning rate and may lead to instabilities. Thus, symexp function makes it faster to reach target values in log space.
@@ D.2 CSCG
CSCG baseline was implemented using code from the repository accompanying the paper (https://github.com/vicariousinc/naturecomm_cscg). CSCG with 10 clones per observation state was trained on batches with a size of 2500 observation steps for Gridworld tasks with 1000 EM algorithm iterations. We iteratively calculated the exponential moving average of transition matrices obtained for different batches with a smoothing coefficient $\alpha=0.9$ as described in the paper [Dedieu et al., 2019]. This smoothed transition matrix was used as an initialization for the next training batch and for inference. The first batch of observations is gathered using the uniform policy, and later batches are gathered using the current best policy. In so-called oracle mode, after the environment changes, the agent gathers a uniform policy batch again.
D.3 DHTM
@@
In the POMDP Gridworld task, a graph for DHTM consists of three hidden variables $H$ with 40 cells per feature state, which is raw observation with 7 states in the case of Gridworld. Each hidden variable gets the same observation, and their predictions are averaged. Using several hidden variables here is essential, as it significantly increases the amount of trajectories memory is able to store by decreasing collision probability for hidden state representations of different trajectories. The same graph is used for the AnimalAI task, but with the encoder block forming one feature variable with 50 states.
In MDP Gridworld, a graph with one hidden variable and one cell for observation state is used, as there is no need to distinguish observations in different contexts.
Generally, the softmax strategy much better fits DHTM than Epsilon greedy. This is connected with the hidden state generalization problem discussed in Section 5. Each new trajectory creates a new hidden state, for which possible future trajectories are not known. The more exploration there is, the less useful the memory is. Softmax renders exploration more narrowed around already-known good trajectories, allowing better reuse of DHTM’s knowledge.
SFs are formed using $\gamma=0.8$ in all experiments except MDP Gridworld, where $\gamma=0.99$ is used. The maximum planning horizon is set to 50 in POMDP Gridworld and AnimalAI, with stopping criteria based on predicted feature variable distribution: 1) DKL between prediction and uniform distribution is less than 0.1, 2) net probability of feature states with positive rewards is greater than 0.01. For MDP Gridworld, planning steps are fixed and equal to 10. The same SF parameters were used for CSCG and LSTM.
In all experiments, segments are initialized with the maximum value, which is $1$. Synapses are initialized with $0.5$. Segments’ and synapses’ learning rates are $0.1$. We also prune segments with values less than $0.3$ to speed up learning.
@@ E.1 Gridworld
We use a simple homebrew Gridworld implementation. Each Gridworld position maps to a colour, reward, and terminal state indicator. There are two special colours for the world border and obstacles. Each timestep, an agent gets the colour index of the current agent’s position. In MDP mode, the agent gets a tuple of its coordinates in the maze. An agent is able to see obstacles or border colour when trying to go through them. We also punish an agent trying to go through a wall or world border in addition to base step reward punishment to facilitate faster convergence to optimal trajectories.
Setups used in our experiments are depicted in Figures 7 and 8. Here, the red colour corresponds to obstacles and the light gray colour to the reward. In each episode, the agent and the terminal rewarding state have the same initial positions. The episode ends when the terminal state is entered or the limit of action steps is reached.
Figure 7: Schematic view of MDP Gridworld setups, in which agent gets its Gridworld position instead of colour index. The left picture shows the initial setup of the environment, with the reward in the center. The right picture represents the environment after the reward is blocked by obstacles (red squares). Agent’s initial position in the lower left corner.
@@ E.2 AnimalAI
AnimalAI is a testbed inspired by experiments with animals [Crosby et al., 2020]. The environment consists of 3D area surrounded by a wall and many different objects that can be placed using a configuration file, including walls, food, ramps, trees, movable obstacles, and more. We tested our agent in a 10x10-meter room surrounded by walls (see Fig. 9). Each timestep, the agent gets an RGB image of a first-person view of the environment and is able to influence its behaviour by choosing one of three actions: turn left, turn right, or go forward. The agent and food item have the same fixed initial positions for each episode. The episode ends if the food item is collected or the maximum episode time is reached. To make an agent’s actions more discretized, we repeat its choice three times and skip three frames.
@@
Appendix F Glossary

### v3

@@ 5 Conclusion
This paper introduces a novel episodic memory model, DHTM. It combines ideas from belief propagation, episodic control, and neurophysiological models of the neocortex, resulting in a fast probabilistic sequence memory.
The experiments show that our memory model can quickly learn observation sequences in changing environments and effectively reuse them for RL tasks, on par with model-based table episodic control. We also show that, unlike CSCG and table EC, DHTM is able to handle distributed representations and account for statistical dependencies between variables of these representations. Potentially, it extends the application of our algorithm to higher state spaces, exploiting their factorizations for more efficient computations and to reduce the number of stored parameters.
Our results suggest that DHTM can be viewed as an adaptive trajectory buffer for a slower generalizing algorithm, such as a Neural State Space Model-based [47]. Through prediction error learning, DHTM can selectively discard and replace obsolete transitions in observed trajectories without disrupting relevant experience. It can also provide an initial policy for an agent through episodic control until a slower but more accurate world model takes over. These features of a trajectory buffer are highly desirable in the case of online learning in non-stationary environments.
Further improvements of DHTM can be directed towards the implementation of an automatic feature space factorization discovery to determine an optimal context factor size during learning. To avoid excessive segment growth in noisy environments, it’s also possible to implement an adaptive segment activation threshold. One of the main limitations of this study is that the evaluation of the algorithms is limited to basic RL setups due to the simple agent architecture. More advanced use cases could be demonstrated using better exploration strategies and in combination with other memory models, which is also a matter of future research.
Appendix A Value Function Decomposition
In our agent model, we approximate the reward function R(s) as a sum:
$R_{t}=\frac{1}{n}\sum^{n}_{k=1}{r(\varphi^{k}_{t})},$ (13)
@@ Appendix A Value Function Decomposition
where $r(\varphi^{k}_{t})$ is a reward associated with state $\varphi^{k}_{t}$, $n$–number of feature variables. Then, similarly to the Successor Representation idea (see Section 2.2), the value function can be represented as:
@@
$=\frac{1}{n}\sum^{n}_{k=1}\sum_{j}\sum^{\infty}_{l=0}\gamma^{l}p(\varphi^{k}_{t+l+1}=j\ |\ h_{t})r^{k}_{j}$
$=\frac{1}{n}\sum^{n}_{k=1}\sum_{j}\mathrm{SF}^{k}_{j}(h_{t})r^{k}_{j},$ (14)
where $\mathrm{SF}^{k}_{j}(h_{t})=\sum^{\infty}_{l=0}\gamma^{l}p(\varphi^{k}_{t+l+1}=j\ |\ h_{t})$, $h_{t}=(h^{1}_{t},...,h^{n}_{t})$—hidden state vector of variables $\{H^{k}_{t}\}_{k}$.
Appendix B Biological Interpretation
@@
The neuronal activity at timestep $t$ can be thought to carry messages $M^{k}_{t-1}$. Messages are propagated through synapses of dendritic segments, which correspond to factors $F^{k}_{c}$. Dendritic segments of biological neurons are known to be coincidence detectors of their synaptic input [45]. We use the notion of dendritic segment to sparsely represent context factors $F^{k}_{c}$, as each factor value corresponds to a particular combination of states (or active cells).
Feature variables $\Phi^{k}_{t}$ may be considered to represent cells of a granular layer (L4), as they are known to be the main hub for sensory excitation for L2/3. L2/3 cells that have common sensory input from the layer L4 are modelled as columns for particular feature states $\mathrm{col(\varphi^{k}_{t})}$ [49].
@@
Appendix C Encoding Observations
Since DHTM is designed to process categorical random variables, we need a categorical encoder for RGB images to solve AnimalAI tasks.
[…]
@@ Appendix D Implementation details
As mentioned in Section 3.4, we incorporate DHTM as a part of an RL agent, which has a memory model and a feature reward function. All tested memory models share the same pipeline—conditioned on actions, they learn sequences of categorically encoded observations and are used to form SFs.
In all experiments, the agent uses softmax exploration with temperature equal to $0.04$, except for EC agent we set softmax temperature to $0.01$, since for higher temperatures its policy is unstable. For SF formation the agent uses $\gamma=0.8$ and $\gamma=0.9$ for Gridworld and AnimalAI experiments respectively. Maximum SF horizon is set to $50$, but it is usually shorter since we employ early stop condition based on feature variables distribution. That is, if the prediction is close to uniform or the probability of a state with associated positive reward meets a threshold, the planning is interrupted. We use probability thresholds equal to $0.05$ and $0.01$ for Gridworld and AnimalAI respectively.
@@ D.1 LSTM
LSTM baseline was implemented with a single LSTMCell from PyTorch library [52]. It is supported by an additional symexp-layer to encode input before passing it to the LSTM cell and a symexp-layer to decode the LSTM cell’s output from the LSTM’s hidden state back to the input representation, where symexp activation function, $\mathrm{symexp}(x)=\mathrm{sign}(x)e^{\lvert x\rvert-1}$, is a reverse of symlog function: $\mathrm{symlog}=\mathrm{sign}(x)\log{(\lvert x\rvert+1})$.
For both Gridworld and AnimalAI tasks, LSTM with a hidden state size of 100 is trained on a buffer of $1000$ trajectories with learning rate $0.002$ on $50$ randomly sampled batches of size $50$ every $10$ episodes.
We also incorporated some notion of random variables in hidden states by splitting the hidden state of the tested LSTM into groups. In all experiments the hidden state represents $10$ categorical variables with $10$ states. That is, LSTM is forced to learn $10$ categorical distributions with multi-cross-entropy loss to explain the observed sequences, which is a somewhat close to the multi-categorical hidden state representation used in DreamerV2/V3 [40]. The idea of using symexp activation function, mentioned above, is inspired by Dreamer too, and is used to remedy the problem of learning extreme probability values. Without symexp the neural network has to represent zero probability with high negative logit values and one-probability with high positive logit values, which is hard to reach with low learning rate and may lead to instabilities. Thus, symexp function makes it faster to reach target values in log space.
@@ D.2 CSCG
CSCG baseline was implemented using code from the repository accompanying the paper (https://github.com/vicariousinc/naturecomm_cscg). CSCG was trained on buffers of $500$ action steps with $1000$ EM steps until convergence. We iteratively calculated the exponential moving average of transition matrices obtained for different batches with a smoothing coefficient $\alpha=0.05$ as described in the paper [53]. This smoothed transition matrix was used as an initialization for the next training batch and for inference. The first batch of observations is gathered using the uniform policy, and later batches are gathered using the current best policy.
In so-called oracle mode, the trajectory buffer is continuously growing, and the transition matrix is trained from scratch every $500$ steps. The buffer is discarded at the environmental change and the transition matrix is reinitialised.
D.3 Episodic Control
Here, we provide an algorithm for SF formation using EC dictionary $\mathcal{D}_{a}$ (see Algorithm 3). In contrast to other memory algorithms, EC dictionary doesn’t estimate state transition probabilities, but it stores connections between them in all-or-nothing faction. Therefore, on each time step, all predicted observations are assumed equally probable. But it’s not a limitation in our set of experiments, since we tested our algorithms in deterministic environments.
Algorithm 3 SF formation for EC
0:  $h_{0}$, $\gamma\in(0,1)$, $T$
0:  $\mathrm{SF}$
1:  $\mathrm{SF}$leftarrow array of zeros
2:  $\mathrm{PS}$leftarrow {$h_{0}$} # $\mathrm{PS}$ is the set of all states (nodes) on the current BFS depth
3:  goal_found $\leftarrow$ false
4:  for $l=1..T$ do
5:     $\mathrm{PS}$leftarrow $\bigcup_{a\in A}\{\mathcal{D}_{a}(h)\}_{h\in\mathrm{PS}}$ get next depth nodes assuming the policy is uniform
6:     counts $\leftarrow$ array of zeros
7:     for all $h\in\mathrm{PS}$ do
[…]
9:        countsφ = countsφ + 1
10:        goal_found $\leftarrow$ $r_{\varphi}>0$ check if the feature state is rewarding
11:     end for
12:     $\mathrm{SF}$ = $\mathrm{SF}$ + $\gamma^{l-1}$ NORMALIZE(counts)
13:     if goal_found is true then
14:        break
15:     end if
16:  end for
D.4 DHTM
@@
In the Gridworld task and AnimalAI with K-Means encoder, the factor-graph for DHTM consists of three hidden variables $H$ with $40$ states per one feature state. Each hidden variable gets the same copied observation, and their predictions are averaged. Using several hidden variables here is essential, as it significantly increases the amount of trajectories memory is able to store by decreasing collision probability for hidden state representations of different trajectories.
In AnimalAI experiments, hidden variables state spaces are of size $40$ states per one feature state. For Categorical VAE encoder, we use one hidden variable per feature variable, except the case (factor 1*) that tests that information from one feature variable is not sufficient for correct prediction, in which three hidden variables per observation variable are used and the factor of size three connecting to only hidden copies of the same feature variable. Examples of typical DHTM’s graphical models used in our experiments are presented on Figure 5. It’s important to note that edge colours denote separate graphical models, but the variables are shared (copied) between them. So there are no loops in these models. For simplicity, we define context factor size as the number of hidden variables from the previous time-step it has connections to, but it also connects to one hidden variable of the current step and the action variable.
[…]
Figure 5: Examples of factor-graphs for two main setups. Edge colours show that edges correspond to separate graphical models with copied variables. All context factors are also connected to action variable $A_{t-1}$. A. Graphical model for one feature variable. In that case, feature value is copied for three hidden variables, to increase hidden state space and avoid memory collisions. B. Graphical model for Categorical VAE distributed encoding, example for factor of size $3$ (i.e. takes into account tree hidden variables from the previous time-step).
@@ E.1 Gridworld
We use a simple homebrew Gridworld implementation. Each Gridworld position maps to a colour, reward, and terminal state indicator. There are two special colours for the world border and obstacles. Each time-step, an agent gets the colour index of the current agent’s position. An agent is able to see obstacles or border colour when trying to go through them, which just replace current position colour for one step without actually changing the agent’s position. We also punish an agent trying to go through a wall or world border in addition to base step reward punishment to facilitate faster convergence to optimal trajectories.
Samples of setups used in our experiments are depicted in Figure 6. Here, the red colour corresponds to obstacles and the light gray colour to the reward. Floor colours are resampled for each trial, but we ensure that there are no more than five floor colours, including the terminal state. In each episode, the agent and the terminal rewarding state have the same initial positions. The episode ends when the terminal state is entered or the limit of action steps is reached.
@@ E.2 AnimalAI
AnimalAI is a testbed inspired by experiments with animals [38]. The environment consists of 3D area surrounded by a wall and many different objects that can be placed using a configuration file, including walls, food, ramps, trees, movable obstacles, and more. We tested our agent in a 10x10-meter room surrounded by walls (see Fig. 7). Each timestep, the agent gets an RGB image of a first-person view of the environment and is able to influence its behaviour by choosing one of three actions: turn left, turn right, or go forward. The agent and food item have the same fixed initial positions for each episode. The episode ends if the food item is collected or the maximum episode time is reached. To make an agent’s actions more discretized, we repeat its choice three times and skip three frames.
@@
Appendix F Additional Experiments
In order to ground our experiments to state-of-the-art model based algorithms, as an additional baseline, we test the smallest version of Dreamer V3 with 18 million parameters. We use the implementation described in Voudouris et al. [54], which is based on Hafner et al. [40] and adapted for AnimalAI.
[…]
Figure 8: Dreamer V3 (18M) score in AnimalAI experiment described in this paper.
Appendix G Glossary